OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 9?

OpenStudy (anonymous):

The possible answers; x2 + y2 = 3 x2 + y2 = 9 x2 + y2 = 81

hartnn (hartnn):

equation of a circle whose center is at the origin and whose radius is r, will be \(\large x^2+y^2=r^2\) here, just plug in r=radius = 9

OpenStudy (anonymous):

81?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^